# Generating CAD Code with Vision-Language Models for 3D Designs

### v1

@@ 4 CADPrompt dataset
For evaluating CADCodeVerify on CAD code generation, we introduce a new benchmark CADPrompt, which consists of $200$ 3D objects, represented both in images and in the Standard Triangle Language (STL) formats. We hired annotators to write natural language prompts for creating each of the objects step-by-step (see examples in Figure 2), and we also hired a CAD design expert to write Python code.
@@ 4.1 Prompt Creation
To construct CADPrompt, we first selected $200$ 3D objects from a dataset of modular CAD objects used in the prior work  (Wu et al., 2021). We then manually annotated each of the $200$ objects with natural language prompts. If a 3D object was abstract or difficult to describe, two annotators would write separate prompts, followed by a third independent annotator who selected the better option. This dual annotation process was conducted for $65$ out of the $200$ examples. Finally, an independent reviewer, who was not involved in the original annotation process, verified and further refined each of the $200$ language prompts to ensure accuracy and grammatical correctness. We chose to exclude the precise numeric parameters (e.g., height of an object) in the prompt as average users may not have access to such knowledge or lack the expertise necessary to specify accurate parametric designs in the real-world use cases. Table 1 provides the statistics of the CADPrompt dataset.
@@ 4.2 Code Annotation
To annotate CADPrompt with Python code, we hired a CAD design expert to develop Python scripts that generate the 3D objects described in §4.1. The expert was provided thelanguage description of the 3D object, the 3D object in STL format, and the geometric properties generated by the geometric solver (see Figure 4). We then use Blender (Flavell, 2011) to verify the 3D objects produced by the Python code with the ground truth, as well as evaluate its geometric properties (see Figure 14). In cases where discrepancies were found, the Python code was returned to the CAD expert for further refinement. Figures 12 and 13 provide examples of Python code from CADPrompt.
@@ 4.3 Data Stratification
We stratify CADPrompt examples according to compilation difficulty and object complexity to gain insights into model performance (§6).
Object complexity:
The complexity of a 3D object is defined by the total number of faces and vertices it contains. The ground truth of the 3D object is represented in mesh file formats, which consist of a collection of vertices, edges, and faces $(x,y,z)$ used to create the 3D object in the STL format. A 3D object with a large number of faces and vertices is considered more complex, while an object with fewer faces and vertices is regarded as simple (see Table 1). We split our dataset by complexity, classifying each 3D object as either (i) Simple or (ii) Complex based on its complexity in relation to the median. Simple objects are defined as those with less complexity than the median while complex objects are greater than the median.
@@
The difficulty of the 3D object is measured by the total amount of compiled code for each object generated during the initial generation step across all experiments. We compare the compiled code resulting from zero-shot and few-shot prompting techniques across all language models utilized in this paper: GPT-4, Genimi, and CodeLlama. We split the data based on the total number of compilable outputs from the experiments (ranging from 0 to 6). Those 3D objects with fewer than four compilable outputs are classified as (i) Easy, whereas 3D objects with $4$ or more compilable outputs are classified as (ii) Hard (see Appendix Figure 11).
@@ 5.1 Baselines
We compare our approach, CADCodeVerify, with two baselines for feedback generation: (i) 3D-Premise, and (ii) Geometric solver feedback.
@@
3D-Premise facilitates code refinement by providing GPT-4 with the image of a generated object, paired with its original description, and prompted GPT-4 to correct any discrepancies between the original and generated objects. To create a baseline against which to we provide the VLM with an image of the generated object and we utilize the same prompts from the original work. Note that this baseline is not applicable for approaches without multimodal capabilities, such as CodeLlama.
@@
We develop a novel baseline that utilizes FreeCAD, an open-source geometric solver, to provide information regarding the geometric dimensions and structure of the generated 3D object, $d_{g}$. However, our proposed baseline, which calculates geometric properties feedback using the geometric solver, requires access to the ground truth 3D objects, $d_{gt}$ in STL format. Specifically, the feedback, $F_{\mathcal{GS}}$, corresponds to numerical parameters across thirteen categories, $S$, each describing a unique geometric dimension or component of the object, e.g., width, height, number of faces, number of vertices, volume. Formally, this feedback process is shown in Equation 6 and 7. We compute $F_{\mathcal{GS}}$ for both the generated design, $d_{g}$, and ground truth, $d_{gt}$, to provide feedback, we can describe the geometric solver as follows:
@@
The feedback from the geometric solver is a numerical value across thirteen categories of geometric information. An example of this feedback, $F_{\mathcal{GS}}$, is shown in Appendix (see Figure 4).This feedback method serves as an upper limit for CAD code refinement, as it conveys the precise geometric differences between the generated 3D object and the ground truth.
We evaluate the performance of each baseline in both zero-shot and few-shot settings. In the few-shot setting, we provide five demonstration examples, with each example consisting of a 3D object description followed by the corresponding code. We use the OpenAI for GPT-4, Replicate API for CodeLlama, and Google’s GenerativeAI API for Gemini 1.5 Pro respectively. [^3] Experiments were conducted from 06 JAN to 15 FEB, 2024, and May 15 to August 15, 2024.
@@ 5.2 Evaluation Metrics
This paper aims to evaluate how accurately VLMs generate and refine 3D objects through CAD code. To evaluate the quality of the generated 3D object compared to the ground truth, we utilize three evaluation metrics: (i) Point Cloud distance, (ii) Hausdorff distance, and (iii) Intersection over the Ground Truth (IoGT). We apply the Iterative Closest Point (ICP) algorithm to rotate and translate the generated 3D object to choose the best possible alignment with the ground truth object (Besl & McKay, 1992). Finally, we normalize each point cloud such that it fits within a unit cube, as per prior work (Zheng et al., 2023). In some cases, the generated CAD code does not compile, which prevents the rendering of a valid 3D object. We report the percentage of successfully compiled code as the “success rate.” The formulas for the evaluation metrics are described as follows:
@@
Point Cloud distance $D(P,Q)$ is shown in Equation 8, where $d_{p,q}=\|p-q\|_{2}$.
@@
Hausdorff distance, $H(P,Q)$, is given by Equation 9, where $\sup$ and $\inf$ represent the supremum and infimum operators, respectively (Beauchemin et al., 1998).
@@
This metric used to evaluate how well the generated 3D object $P$ aligns with the ground truth $Q$. As shown in Equation 10, it calculates the area of intersection between $P$ and $Q$ to the area of $Q$.
@@
Point Cloud distance and Hausdorff distance are based on the geometric properties (e.g., height, width, and volume) of 3D objects, while the IoGT measures the overlap of the bounding boxes between the generated and ground truth 3D objects. When computing these distances, we assign the maximum distance of two points within a unit cube, i.e., $\sqrt{3}$, and an IoGT value of zero to any data point where the code fails to compile. This approach maximally penalizes unsuccessful 3D object generations, ensuring consistent and fair evaluation criteria across all the experiments.
Table 2: We present our results using the median Point Cloud distance, Hausdorff distance, and Intersection over Ground Truth (IoGT), along with the interquartile range (IQR) and success rates. The * symbol indicates that the geometric solver accesses the ground truth to compute the geometric differences between the ground truth and the generated 3D object.
Model Feedback Mechanism IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Success Rate $\uparrow$
@@ 6 Experiment Results
We utilize CADPrompt to quantitatively evaluate the capabilities of CADCodeVerify approach across various VLMs. The following are our key findings:
@@
GPT-4 exhibits the highest capacity to generate compile-able code on CADPrompt. We report our primary results in Table 2, comparing the distance and success-rate for three LLMs and three refinement methods. With regards to “success-rate,” GPT-4 exhibits superior performance $96.5\%$ on the CADPrompt benchmark compared to Gemini $85\%$ and CodeLlama $73.5\%$. While prior work has provided valuable qualitative insights into GPT-4’s competency to producing compile-able code (Makatura et al., 2023), our analysis, via CADPrompt, provides a concrete assessment of each LLMs CAD code generation capacity.
@@
CADCodeVerify shows superior performance on more challenging data. Figure 3 presents the success rates for each feedback mechanism at generation and refinement stages, using the difficulty and complexity splits of CADPrompt. For "Easy" data, the performance across all baselines is relatively comparable. However, results on the "Hard" data underscore the effectiveness of CADCodeVerify. It is the only refinement approach that improves the success rate of object generation, achieving approximately a 9% increase at the Refine-1 stage. In contrast, 3D-Premise struggles to provide effective feedback for "Hard" data, resulting in a 20% drop in success rate at Refine-1, bringing it down to 62%. Real-world data in design and manufacturing is likely to align more closely with the "Complex" and "Hard" data splits. Our results demonstrate the value of CADCodeVerify in providing feedback that can be applied to real-world product design.
@@
Figure 3: Success rate comparison across different feedback mechanisms, categorized by difficulty and complexity. Each figure corresponds to a specific subset of the CADPrompt dataset, covering both generated and refined objects. The plots for the “Hard” and “Complex” splits indicate that the feedback approach from 3D-Premise significantly reduces GPT-4’s success rate in object generation.
@@
CADCodeVerify generates model-agnostic feedback to improve 3D object generation. Table 2 also reports the distance measure from the ground truth at the generate and refine stages. Since CodeLlama does not possess multimodal capabilities, we utilize GPT-4 to perform CADCodeVerify to generate refinement feedback. Our results show that CADCodeVerify improves the quality of the generated object across all three LLMs/VLMs as measured by IoGT, Point Cloud distance and Hausdorff distance. CADCodeVerify also improves both GPT-4 and CodeLlama’s “success-rate” in producing compile-able CAD code. This result emphasizes the model-agnostic nature of the multimodal “assess and reason” analysis induced by our CADCodeVerify refinement process, to universally produce actionable feedback for correcting object generation errors.
Table 3: The main results and ablation analysis for the GPT-4 few-shot setting on a subset of 100 3D objects. Results are reported using the median and interquartile range (IQR).
@@
Ablation Study IoGT $\uparrow$ PLC distances $\downarrow$ Hausdorff dist. $\downarrow$ Success Rate $\uparrow$
@@
CADCodeVerify outperforms 3D-Premise across VLMs. In Table 2, we quantitatively compare our CADCodeVerify approach to a refinement approach proposed in prior work, 3D-Premise. Since 3D-Premise involves directly uploading an image of the generated object to the VLM, this approach cannot be applied to the unimodal CodeLlama model. With respect to GPT-4 and Gemini, CADCodeVerify is measured to more accurately refine objects, compared to 3D-Premise for objects in the CADPrompt dataset ($>7\%$ reduction in point cloud distance for GPT-4 and Gemini) . This result shows that CADCodeVerify is the new state-of-the-art for CAD code refinement. Next, we compare our proposed refinement approach to an upper benchmark which utilizes a geometric solver to provide parametric feedback regarding the specific, geometric differences with regards to the target object, as per § 5. Our results show that through CADCodeVerify, we can achieve comparable performance without needing to access the target object, which is unlikely to be available in real-world use-cases.
@@
Ablation Analysis. We perform an ablation analysis to test the effectiveness of each independent component of CADCodeVerify (See Table 2). We conduct this ablation with GPT-4, with a few-shot prompt, on 100 randomly selected examples from CADPrompt. First, we ablate the inclusion of few-shot example questions in the question-generation component (see Equation 3). The results shows that zero-shot QA generation increases the distance between the ground truth and refined 3D objects from $0.126$ to $0.141$ in Point Cloud distance and from $0.444$ to $0.471$ in Hausdorff distance. Our second ablation measures the impact of the reference images, $I_{ref}$, during the code refinement step of Equation 3. Removing the reference images worsens the Point Cloud distance and Hausdorff distances from $0.126$, $0.444$ to $0.153$, $0.451$ respectively, highlighting the importance of the reference images.

### v2

@@ 4 CADPrompt dataset
For evaluating CADCodeVerify on CAD code generation, we introduce a new benchmark the CADPrompt, which consists of $200$ 3D objects, represented both in images and STL. Each object in CADPrompt is annotated with ground truth code, and a natural language prompt.
@@ 4.1 Prompt Creation
To construct CADPrompt dataset, we first selected $200$ 3D objects from a collection of modular CAD objects from previous work  (Wu et al., 2021). Each object was manually annotated with a natural language prompt. For difficult-to-describe objects, two annotators independently provided prompts. An independent third annotator then selected the more suitable prompt. Finally, a fourth independent reviewer, not involved in the original annotation, verified and refined each of the $200$ prompts to ensure accuracy and grammatical correctness. Table 1 provides the statistics of CADPrompt.
@@ 4.2 Code Annotation
We recruited a CAD Design expert to annotate CADPrompt with ground truth code for each object. The CAD design expert was provided with the language description of the 3D object, the object in STL format, and its geometric properties generated by the geometric solver (Figure 6). We then used Blender (Flavell, 2011) to validate the 3D objects produced by the expert’s Python code against the ground truth and to evaluate their geometric properties (Figure 18). In cases where discrepancies were identified, the Python code was returned to the CAD expert for additional refinement. Figures 16 and 17 provide examples of Python code from CADPrompt.
@@ 4.3 Data Stratification
We stratify CADPrompt examples by mesh complexity, geometric complexity and compilation difficulty to gain insights into model performance (§6).
Mesh complexity:
We define “mesh complexity” as the total number of faces and vertices of an object. The ground truth of the 3D object is represented in mesh file formats, which consist of a collection of vertices, edges, and faces $(x,y,z)$ used to create the 3D object in the STL format. Objects with more faces and vertices are classified as more complex, while those with fewer faces and vertices are considered simple (Table 1). We split the dataset into two groups based on the median complexity: (i) Simple (those with fewer faces and vertices than the median) and (ii) Complex objects (those with more).
@@
We define “compilation difficulty” to be a measure of how difficult it is for a set of three language models (i.e., GPT-4, Gemini, and CodeLlama) to generate code for a given 3D object across two prompting methods (i.e., zero- and few-shot prompting), for a total of six attempts to generate compilable code. 3D objects were labeled then as either (i) Easy (at least four of six methods generated compilable code) and (ii) Hard (otherwise). See Appendix Figure 15.
Figure 3: Examples of 3D objects from the CADPrompt dataset.
Geometric complexity:
We enlisted a CAD design expert to evaluate the “geometric complexity” of each object in CADPrompt. Each object was assigned one of the following levels: (i) Simple: the object is basic, with few features. It may consist of one geometric shape; (ii) Moderate: the object has a moderate amount of detail, with a few distinct features or components; (iii) Complex: the object has many interconnected parts, fine details, or intricate shapes; and (iv) Very Complex: the object is highly intricate, with many components, detailed textures, or complex shapes. It may have a large number of fine details, interlocking parts, or unique geometric features (Figure 3).
@@ 5.1 Baselines
We compare our approach, CADCodeVerify, with two baseline methods for feedback generation: (i) 3D-Premise, and (ii) Geometric solver feedback.
@@
3D-Premise facilitates code refinement by providing GPT-4 with both the image of a generated object and its original description. GPT-4 is then prompted to correct any discrepancies between the original description and generated objects. To establish a baseline for comparison, we supply the VLM with an image of the generated object and use the same prompts as used in the original work. Note that this baseline is not applicable for approaches lacking multimodal capabilities, such as CodeLlama.
@@
We develop a novel baseline that leverages FreeCAD, an open-source geometric solver, to provide information about the geometric dimensions and structure of the generated 3D object, $d_{g}$. However, our proposed baseline, which computes geometric properties feedback using the geometric solver, requires access to the ground truth 3D objects, $d_{gt}$ in STL format. Specifically, the feedback, $F_{\mathcal{GS}}$, consists of numerical values across thirteen categories, $S$, each representing a unique geometric dimension or component of the object, such as width, height, number of faces, number of vertices, and volume. Formally, this feedback process is shown in Eq. 6 and 7. We compute $F_{\mathcal{GS}}$ for both the generated design, $d_{g}$, and ground truth, $d_{gt}$. In summary, we can describe the role of the geometric solver as follows:
@@
The feedback from the geometric solver consists of numerical values across thirteen categories of geometric information. An example of this feedback, $F_{\mathcal{GS}}$, is shown in Appendix (see Figure 6). This feedback method acts as an upper bound for CAD code refinement, as it conveys the exact geometric differences between the generated 3D object and the ground truth.
@@ 5.2 Evaluation Metrics
This evaluates how accurately VLMs generate and refine 3D objects through CAD code. We use three evaluation metrics to compare the quality of the generated 3D object against the ground truth: (i) Point Cloud distance, (ii) Hausdorff distance, and (iii) Intersection over the Ground Truth (IoGT). We apply the Iterative Closest Point (ICP) algorithm to rotate and translate the generated 3D object for optimal alignment with the ground truth object (Besl & McKay, 1992). Finally, each point cloud is normalized to fit within a unit cube, as per prior work (Zheng et al., 2023). We also report the percentage of successfully compiled code as the “compile rate”. The formulas for the evaluation metrics are described as follows:
@@
Point Cloud distance $D(P,Q)$ is shown in Eq. 8, where $d_{p,q}=\|p-q\|_{2}$.
@@
Hausdorff distance, $H(P,Q)$, is given by Eq. 9, where $\sup$ and $\inf$ represent the supremum and infimum operators, respectively (Beauchemin et al., 1998).
@@
This metric used to evaluate how closely the generated 3D object $P$ aligns with the ground truth $Q$. As shown in Eq. 10, it calculates the ratio of the intersection area between $P$ and $Q$ to the area of $Q$.
@@
Point Cloud distance and Hausdorff distance are based on the geometric properties (e.g., height, width, and volume) of 3D objects, while the IoGT measures the overlap of the bounding boxes between the generated and ground truth 3D objects. If any generation fails to compile, we set the distance (Point Cloud and Hausdroff) to $\sqrt{3}$ (the largest possible distance between two points in a unit cube) and an IoGT value of zero (the worse possible IoGT value between two 3D objects) to maximally and uniformly penalize unsuccessful 3D object generations.
Table 2: This table reports the median (IQR) benchmarking results for baselines across metrics. The * symbol indicates that the geometric solver accesses the ground truth to compute the geometric differences between the ground truth and the generated 3D object.
Model Feedback Mechanism IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Compile Rate $\uparrow$
@@ 6 Experiment Results
We utilize CADPrompt to quantitatively evaluate the capabilities of CADCodeVerify approach across various VLMs. In the rest of this section, we refer to the three stages of the CAD code generation process as “Generate” “Refine-1” and “Refine-2” “Generated” refers to the object generated by the LLM after the Code-Execution Step. “Refine-1” refers to the object after the first step of refinement, and Refine-2 refers to the object after the second step of refinement. Our key findings are as follows:
@@
GPT-4 demonstrates the highest capacity to generate compilable code CADPrompt. We present our primary results in Table 2, comparing the distance and success-rate across three LLMs and three refinement methods. Regarding compile rate, GPT-4 shows superior performance at 96.5% on the CADPrompt benchmark, compared to Gemini at 85% and CodeLlama at 73.5%. While previous studies have provided valuable qualitative insights into GPT-4’s ability to produce compilable code (Makatura et al., 2023), our analysis using CADPrompt offers a concrete evaluation of each LLM’s CAD code generation capability.
@@
CADCodeVerify demonstrates superior performance on more challenging data. Figure 4 shows the compile rates for each feedback mechanism at the generation and refinement stages, using the difficulty and complexity splits of CADPrompt. For "Easy" data, the performance across all baselines is relatively similar. However, results on the "Hard" data underscore the effectiveness of CADCodeVerify, which is the only refinement approach that enhances the compile rate of object generation, achieving approximately a 9% increase at the Refine-1 stage. In contrast, 3D-Premise struggles to provide effective feedback for "Hard" data, resulting in a 20% drop in compile rate at Refine-1, bringing it down to 62%. Real-world data in design and manufacturing is likely to align more closely with the “Complex” and “Hard” data splits, underscoring the value of CADCode. Our findings underscore the value of CADCodeVerify in delivering feedback that can be applied to real-world product design.
@@
Figure 4: Compile rate comparison across different feedback mechanisms, categorized by difficulty and mesh complexity. Each figure corresponds to a specific subset of the CADPrompt dataset, covering both generated and refined objects. The plots for the “Hard” and “Complex” splits indicate that the feedback approach from 3D-Premise reduces GPT-4’s compile rate in object generation.
@@
CADCodeVerify generates model-agnostic feedback to improve 3D object generation. Table 2 presents the distance measure from the ground truth at the generation and refinement stages. Since CodeLlama lacks multimodal capabilities, we employ GPT-4 to execute CADCodeVerify and generate refinement feedback. Our results indicate that CADCodeVerify improves the quality of the generated object across all three LLMs/VLMs as measured by IoGT, Point Cloud distance and Hausdorff distance. CADCodeVerify also increases both GPT-4 and CodeLlama’s “compile rate” in producing compilable CAD code. This result highlights the model-agnostic nature of the multimodal “assess and reason” analysis induced by our CADCodeVerify refinement process, to universally provides actionable feedback for correcting object generation errors.
Table 3: Median (IQR) results from our ablation study conducted on a randomly selected subset of 100 samples.
@@
Ablation Study IoGT $\uparrow$ PLC distances $\downarrow$ Hausdorff dist. $\downarrow$ Compile Rate $\uparrow$
@@
CADCodeVerify outperforms 3D-Premise across VLMs. In Table 2, we quantitatively compare our CADCodeVerify approach to a refinement method proposed in prior work, 3D-Premise. Since 3D-Premise requires directly uploading an image of the generated object to the VLM, this approach cannot be applied to the unimodal CodeLlama model. For GPT-4 and Gemini, CADCodeVerify is shown to refine objects more accurately than 3D-Premise for objects in the CADPrompt dataset ($>7\%$ reduction in point cloud distance for GPT-4 and Gemini). This finding establishes CADCodeVerify as the new state-of-the-art for CAD code refinement. Next, we compare our proposed refinement approach to an upper benchmark that uses a geometric solver to provide parametric feedback on specific geometric differences relative to the target object, as per § 5. Our results show that through CADCodeVerify, we can achieve comparable performance without requiring access to the target object, which is typically unavailable in real-world applications.
@@
Ablation analysis. We conduct an ablation to evaluate the effectiveness of each independent component of CADCodeVerify (see Table 2). This analysis is performed using GPT-4 with a few-shot prompt on 100 randomly selected examples from CADPrompt. First, we test the effect of removing the few-shot example questions in the question-generation component (see Eq. 3). The results indicate that zero-shot QA generation increases the distance between the ground truth and refined 3D objects from $0.126$ to $0.141$ in Point Cloud distance and from $0.444$ to $0.471$ in Hausdorff distance. Our second ablation measures the impact of the reference images, $I_{ref}$, during the code refinement step of Eq. 3. Excluding the reference images worsens the Point Cloud and Hausdorff distances from 0.126 and $0.444$ to $0.153$ and $0.451$, respectively, emphasizing the importance of the reference images.
Table 4: Performance Comparison of CADCodeVerify and Human-in-the-Loop (HITL) approaches on a subset of 50 examples from the GPT-4 few-shot setting.
Model Feedback Mechanism IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Compile Rate $\uparrow$
GPT-4: Few-shot Generated 0.930 (0.043) 0.156 (0.138) 0.495 (0.287) 98.5%
CADCodeVerify 0.948 (0.036) 0.137 (0.136) 0.445 (0.302) 98.5%
Human-in-the-Loop 0.944 (0.032) 0.120 (0.140) 0.397 (0.354) 99.0%
Gold-Standard human feedback outperforms CADCodeVerify slightly. This experiment evaluates the performance of CADCodeVerify compared to a human-in-the-loop approach, where instead of using CADCodeVerify for code refinement, we provide gold standard language feedback which includes the exact changes that need to be made to the object. We performed this experiment on a randomly selected subset of 50 examples using the GPT-4 few-shot setting. The experiment involves two steps: (1) a human participant selects the image with the best viewing angle of the 3D object from four options (see Eq. 3), and (2) the human provides written feedback. The results indicate that the human-in-the-loop approach led to a slight improvement in Point Cloud distance and Hausdorff distance, with performance increasing from $0.137$ and $0.445$ to $0.120$ and $0.397$, respectively (see Table 4). The result of the human-in-the-loop contextualizes the performance of CADCodeVerify with respect to a method which injects domain-expertise to provide the best possible feedback. While CADCodeVerify makes significant strides in refining 3D objects, without any requiring domain expertise, some improvements are still needed to reach the level of a human expert.
Figure 5: An example of a subset of questions and answers generated by CADCodeVerify.
Accuracy of generated answers. In the Question-Answering phase, CADCodeVerify is prompted to respond with “Yes,” “No” or “Unclear,” in situations when there is insufficient information (See Figure 5). To evaluate the accuracy of these answers, we randomly selected a subset of 50 examples from the GPT-4 few-shot setting, then manually validated the answers for both refinement stages. The results indicate that CADCodeVerify provides correct answers with an accuracy of 64.6% for Refine 1 and 68.2% for Refine 2 (see Table 9). To reduce hallucinations from the LLMs, we instructed it to respond with "Unclear" whenever it lacked confidence in its answers. In future work, we aim to explore how LLMs can interpret 3D objects and investigate methods to teach LLMs to self-verify their generated responses.